OpenStudy (anonymous):

HELP, WILL REWARD! Solve the exponential equation. 125^7x-2=150 a.-0.1375 b.2.1483 c.0.4234 d.0.4340

OpenStudy (anonymous):

is it like this? \[125^{(7x-2)}=150\]

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

have you learned about logarithms?

OpenStudy (anonymous):

yes i just dont get this equation

OpenStudy (anonymous):

so take the log (or natural log) of both sides... \[ 125^{(7x-2)}=125\Rightarrow \ln\left(125^{(7x-2)}\right)=\ln125\]\[\Rightarrow (7x-2)\ln125=\ln125 \Rightarrow 7x-2=\frac{\ln125 }{\ln125}\] can you solve it from here?

OpenStudy (anonymous):

oops, i mean like this \[7x-2=\frac{\ln150}{\ln125}\]

OpenStudy (anonymous):

I got 3.74.

OpenStudy (anonymous):

let's look... \[7x-2=\frac{\ln150}{\ln125} \Rightarrow x=\left(2+\frac{\ln150}{\ln125}\right)\div 7=(2+1.0377609175...)\div 7=\frac{3.0377609175...}{7}\]

OpenStudy (anonymous):

this should be less than .5.

OpenStudy (anonymous):

i cant see the full equation that you wrote.

OpenStudy (anonymous):

\[x=\frac{3.077609175...}{7}\]

OpenStudy (anonymous):

0.439 is what i got but its none of the answer choices

OpenStudy (anonymous):

i got .4339... which rounds to

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

are you asking or telling?

OpenStudy (anonymous):

asking?

OpenStudy (anonymous):

k, you should be telling. the whole idea is to get confortable and then confident

OpenStudy (anonymous):

are you unsure of any of the steps?

OpenStudy (anonymous):

Never said i wasnt comfortable or confident so yea. Nothing wrong with asking dude.

OpenStudy (anonymous):

I get it.

OpenStudy (anonymous):

you're right, nothing wrong with asking. don't get me wrong, i just want you to understand how to do these. if you're asking if something is right then my feeling is that you're unsure of something, that's all. and i want to know where that's occurring so i can do my best to help you with that.